# RevPRAG: Revealing Poisoning Attacks in Retrieval-Augmented Generation through LLM Activation Analysis

### v3

@@ 5.3 Ablation Study
Effects of different support set size. In RevPRAG, the Support Data plays a pivotal role in the model’s learning process. It supplies labeled information and task-specific knowledge, allowing the model to conduct effective reasoning and learning even with limited data. We experiment with various support set sizes ranging from 50 to 250 to examine their effect on the performance of RevPRAG. This evaluation was conducted on Llama2-7B with different datasets. The results in Fig. 6 indicate that varying the support size does not significantly impact the model’s detection performance.
@@ 6 Conclusion
Appendix A Additional Results
Table 6: Generalization Performance of RevPRAG for RAG with four different LLMs. The training and test datasets vary across different rows. Abbreviations: Hot (HotpotQA), MS (MS-MARCO).
Training Dataset Test Dataset Metrics LLMs of RAG
GPT2-XL 1.5B Llama2-7B Mistral-7B Llama3-8B
NQ & Hot MS TPR 0.881 0.886 0.948 0.956
FPR 0.134 0.149 0.076 0.066
Hot & MS NQ TPR 0.980 0.983 0.988 0.980
FPR 0.007 0.074 0.078 0.038
NQ & MS Hot TPR 0.977 0.961 0.942 0.978
FPR 0.025 0.089 0.055 0.049
NQ & Hot & MS NQ & Hot & MS TPR 0.986 0.994 0.985 0.987
FPR 0.032 0.007 0.009 0.035
A.1 Generalization
Given the wide range of RAG application scenarios and the diverse user requirements it faces, it is impractical to ensure that our detection model has been trained on all possible scenarios and queries in real-world applications. However, the performance of neural network models largely depends on the similarity between the distributions of the training data and the test data Yang et al. (2024). Consequently, our model’s performance may degrade when faced with training and test data that stem from differing distributions—a challenge frequently observed in real-world scenarios. To address this issue, we conduct a series of generalization experiments. Specifically, we train the detection model using any two datasets and test it on a third dataset that was not used during training. For example, we use NQ and HotpotQA as training datasets and MS-MARCO as the testing dataset. Although these three datasets are all QA datasets, they exhibit significant differences. For example, NQ focuses on extracting answers to factual questions from a single long document, HotpotQA involves multi-document reasoning to derive answers, and MS-MARCO retrieves and ranks relevant answers from a large-scale collection of documents. Therefore, conducting generalization experiments based on these three datasets is reasonable.
Table 6 illustrates the TPRs and FPRs of RevPRAG for RAG with four different LLMs. Overall, the experimental results demonstrate that our detection model exhibits strong generalization performance across RAG with different LLMs and various datasets. For example, when using HotpotQA and MS-MARCO as training data, the detection model achieves TPRs of 98% (with GPT2-XL 1.5B), 98.3% (with Llama2-7B), 98.8% (with Mistral-7B), and 98% (with Llama3-8B) on the NQ dataset. Meanwhile, all FPRs remain below 8%. Furthermore, we observe that the generalization performance is best when NQ is used as the test data (for instance, 98.3% with Llama2-7B), while MS-MARCO shows the poorest performance (for instance, 88.6% with Llama2-7B). We attribute this to the fact that the questions and tasks in HotpotQA and MS-MARCO are more complex compared to those in NQ. Therefore, detection models trained on more complex tasks generalize well to simpler tasks, whereas the reverse is more challenging. In conclusion, these experimental results highlight that RevPRAG exhibits strong generalization and robust detection performance, even in the presence of significant discrepancies between the training and test datasets.
(a) TPRs of RevPRAG.
(b) FPRs of RevPRAG.
Figure 7: RevPRAG trained on the activations from specific layers.
A.2 RevPRAG’s Performance on Complex Open-Ended Questions
In this section, we conducted a series of experiments to evaluate the performance of RevPRAG on complex, open-ended questions (e.g., “how to make relationship last?”). These questions present unique challenges due to their diverse and unstructured nature, in contrast to straightforward, closed-ended questions (e.g., “What is the name of the highest mountain?”). In our experiments, the NQ, HotpotQA, and MS-MARCO datasets primarily consist of close-ended questions. As a result, the majority of our previous experiments focused on close-ended problems, which was our default experimental setting. In this study, we utilized the advanced GPT-4o to filter and extract 3,000 open-ended questions from the HotpotQA and MS-MARCO datasets for training and testing the model. For open-ended questions, cosine similarity is employed to evaluate whether the LLM’s response aligns with the attacker’s target response. If the similarity surpasses a predefined threshold, it is considered indicative of a successful poisoning attack.
Dataset Metrics LLMs of RAG
GPT2-XL 1.5B Llama2-7B Mistral-7B Llama3-8B
HotpotQA TPR 0.982 0.995 0.991 0.982
FPR 0.033 0.029 0.008 0.007
MS-MARCO TPR 0.988 0.989 0.990 0.983
FPR 0.009 0.009 0.001 0.017
Table 8: RevPRAG trained on the activations from specific individual layers of GPT2-XL 1.5B.
Dataset Metrics Different layers
layer 0 layer 8 layer 15 layer 22 layer 29 layer 36 layer 41 layer 47
NQ TPR 0.996 0.988 0.996 0.984 0.996 0.988 0.992 0.996
FPR 0.027 0.007 0.017 0.003 0.007 0.003 0.017 0.003
HotpotQA TPR 0.713 0.984 0.994 0.989 0.520 0.619 0.931 0.992
FPR 0.409 0.023 0.012 0.006 0.445 0.409 0.023 0.019
MS-MARCO TPR 0.967 0.998 0.988 0.986 0.988 0.963 0.955 0.992
FPR 0.023 0.004 0.002 0.019 0.030 0.026 0.037 0.017
A.3 Activations from Specified Layers.
Fig. 7 illustrates the detection performance of RevPRAG using activations from different layers of various LLMs. In previously presented experiments, we utilize activations from all layers as both training and testing data, yielding excellent results. Additionally, we also test using different layers. The experimental results in Fig. 7 demonstrate that utilizing activations from only the first few layers can still achieve satisfactory detection performance, providing valuable insights for future research. For example, when using activations from layers 0 to 5, RevPRAG achieved TPRs exceeding 97% while maintaining FPRs below 7% for RAG with all LLMs on HotpotQA. However, the experimental results also suggest that using activations from intermediate or deeper layers can lead to performance fluctuations, including signs of degradation or slower convergence. For instance, when using activations from layers 16 to 24 with Llama3-8B as the LLM in RAG, RevPRAG achieves a TPR of 78.8% on NQ dataset and 86% on MS-MARCO dataset.
We further explored the use of activations from a specific individual layer of the LLMs to train and test RevPRAG. We chose 8 layers with roughly even spacing for testing. As shown in Table 8, when using activations from only a specific layer of the GPT2-XL model, RevPRAG demonstrates excellent performance in general. For instance, when the model is trained using activations from layer 0 on the NQ dataset, the TPR can reach as high as 99.6%. However, we also observed that activations from certain layers do not yield satisfactory performance. For example, when the model is trained using activations from layer 29 on the HotpotQA dataset, the TPR is only 52%, while the FPR reaches 44.5%. It is precisely due to the existence of these suboptimal layers that models trained with multi-layer activations may not always outperform those using single-layer activations (such as layer 0 with NQ dataset). However, incorporating multi-layer activations can enhance the model’s stability, mitigating the detrimental effects of these suboptimal layers.
A.4 Impact of Similarity Metric.
Table 9 shows the results on the HotpotQA dataset, indicating that the choice of similarity calculation method has minimal impact on RevPRAG’s performance, which consistently achieves high TPR and low FPR. This suggests the robustness of our approach, as it reliably identifies poisoned texts even when LLM activations vary slightly under similar conditions.
Table 9: Impact of similarity metric.
Similarity Metric Metrics LLMs of RAG
GPT2-XL 1.5B Llama2-7B Mistral-7B Llama3-8B
Dot Product TPR 0.972 0.985 0.977 0.973
FPR 0.016 0.061 0.022 0.017
Cosine TPR 0.978 0.990 0.979 0.981
FPR 0.037 0.011 0.023 0.043
A.5 Isolating Poisoned Responses and Hallucinations
It is well-known that hallucinations are an inevitable phenomenon in LLMs. Even with the introduction of a knowledge database in RAG, LLMs may still generate non-factual responses due to hallucinations. Therefore, the incorrect responses generated by RAG may also stem from hallucinations, rather than being solely caused by RAG poisoning. We conducted experiments to test if our approach can distinguish hallucinations and RAG poisoning. Fig. 8 shows the t-SNE representation of mean activations for poisoned response and hallucinations across all layers for Mistral-7B and Llama2-7B on the NQ dataset. We observe that activations across all layers clearly distinguish between hallucinations and poisoned responses.
Figure 8: t-SNE visualizations of activations for poisoned responses and hallucinations.
This key finding has led us to extend our approach to differentiate between poisoned responses and hallucinations. We thus continue to collect data and train the model using the process outlined in Fig. 3, with the only difference being that we now collect hallucination data. We also conduct extensive experiments on RAG with different LLMs and datasets. From the experimental results in Table 10, we can see that our method achieves a high TPR across all LLMs and datasets. For instance, RevPRAG achieved 98.7% (on NQ), 97.5% (on HotpotQA), and 97.3% (on MS-MARCO) TPRs for RAG with GPT2-XL 1.5B. Furthermore, we observe that the FPR remains low across all evaluation settings. As shown in the table, RevPRAG could achieve 0.8% (on NQ), 0.8% (on HotpotQA) and 0.6% (on MS-MARCO) FPRs for RAG with Llama3-8B. This further supports our previous observation that there is a clear distinction between poisoned responses and hallucinations.
Table 10: RevPRAG could achieve high TPRs and low FPRs to distinguish poisoned responses and hallucinations.
Dataset Metrics LLMs of RAG
GPT2-XL 1.5B Llama2-7B Mistral-7B Llama3-8B
NQ TPR 0.987 0.983 0.993 0.995
FPR 0.046 0.017 0.069 0.008
HotpotQA TPR 0.975 0.978 0.991 0.995
FPR 0.004 0.058 0.004 0.008
MS-MARCO TPR 0.973 0.984 0.999 0.989
FPR 0.009 0.023 0.001 0.006

### v4

@@ 5.3 Ablation Study
Effects of different support set size. In RevPRAG, support data provides essential labeled and task-specific information, facilitating effective reasoning and learning under limited data conditions. We experiment with various support set sizes ranging from 50 to 250 to examine their effect on the performance of RevPRAG. The results in Fig. 5 indicate that varying the support size does not significantly impact the model’s detection performance.
@@
5.4 RevPRAG’s Performance on Complex Open-Ended Questions
In this section, we conducted a series of experiments to evaluate the performance of RevPRAG on complex, open-ended questions (e.g., “how to make relationship last?”). These questions present unique challenges due to their diverse and unstructured nature, in contrast to straightforward, closed-ended questions (e.g., “What is the name of the highest mountain?”). In our experiments, the NQ, HotpotQA, and MS-MARCO datasets primarily consist of close-ended questions. As a result, the majority of our previous experiments focused on close-ended problems, which was our default experimental setting. In this study, we utilized the advanced GPT-4o to filter and extract 3,000 open-ended questions from the HotpotQA and MS-MARCO datasets for training and testing the model. For open-ended questions, cosine similarity is employed to evaluate whether the LLM’s response aligns with the attacker’s target response. If the similarity surpasses a predefined threshold, it is considered indicative of a successful attack.
Dataset Metrics LLMs of RAG
GPT2-XL 1.5B Llama2-7B Mistral-7B Llama3-8B
HotpotQA TPR 0.982 0.995 0.991 0.982
FPR 0.033 0.029 0.008 0.007
MS-MARCO TPR 0.988 0.989 0.990 0.983
FPR 0.009 0.009 0.001 0.017
5.5 Effect of Real-world Natural Text Noise on Detection Performance
To better approximate real-world scenarios, we injected natural textual noise (e.g., spelling and grammatical errors) into clean texts using the GPT-4o model. Noise was introduced in a controlled manner to preserve semantics while simulating realistic imperfections. Specifically, 50% of the clean texts were modified, with noise accounting for 10% of the total word count in each case. Notably, noise injection was applied only to clean texts, as poisoned texts are typically well-formed to maximize attack effectiveness and thus unlikely to contain such artifacts.
As shown in Table 7, the experimental results indicate that injecting a certain amount of natural noise into clean texts does not significantly affect the detection performance of our proposed method. This further demonstrates the robustness of our approach. We attribute this to two main factors. First, the textual perturbations introduced by poisoning attacks are fundamentally different in nature from natural noise, and this distinction is clearly reflected in the activation vectors of the LLM. Second, LLMs are inherently robust to natural noise, which does not substantially interfere with their ability to generate correct responses.
Table 7: Performance of RevPRAG on noisy datasets.
Dataset Metrics GPT2-XL 1.5B Llama3-8B
Noisy_NQ TPR 0.980 0.977
FPR 0.034 0.012
Noisy_HotpotQA TPR 0.969 0.974
FPR 0.026 0.011
Noisy_MS-MARCO TPR 0.989 0.977
FPR 0.013 0.018
5.6 Efficiency
Table 8 compares the time overhead between LLM Factoscope He et al. (2024) and RevPRAG when the LLM in RAG is Llama3-8B, including the average training time per epoch and the average inference time per test sample. This experiment was conducted using 1,000 training samples and 500 test samples, with poisoned and clean examples each accounting for 50%. The results demonstrate that RevPRAG, with its task-specific architecture and carefully selected detection metrics, incurs significantly lower computational costs than LLM Factoscope, which integrates multiple sub-models for hallucination detection. Its efficient detection capability makes RevPRAG particularly well-suited for latency-sensitive RAG scenarios, underscoring its practical value.
Table 8: Comparison of time overhead.
Dataset Training Time per Epoch Inference Time per Sample
LLM factoscope RevPRAG LLM factoscope RevPRAG
NQ 91.61s 19.31s 0.0051s 0.0021s
HotpotQA 101.25s 23.69s 0.0066s 0.0023s
MS-MARCO 94.47s 20.72s 0.0058s 0.0022s
@@ 6 Conclusion
Appendix A Training Details
A.1 Dataset
As shown in Table 9, we present the average response lengths for both poisoned and correct answers generated by GPT2-XL across three datasets (NQ, HotpotQA, and MS-MARCO), along with examples illustrating each answer format for a specific question. To evaluate the detection of poisoning attacks on the knowledge base of RAG, we selected 3,000 instances of triples $(q,t,a)$ from each of the three evaluation datasets mentioned above. In each triple, $q$ denotes a question, $t$ represents the supporting text collected from Wikipedia or web documents corresponding to $q$, and $a$ is the correct answer to $q$, generated using the state-of-the-art GPT-4 model. Among these 3,000 triplets, 1,500 are randomly selected as benign instances, while the remaining 1,500 are designated as poisoned instances. For each poisoned instance, the poisoned answer $a_{p}$ is generated by GPT-4 for the given question $q$, and the poisoned text $t_{p}$ is crafted using existing poisoning strategies, including PoisonedRAG Zou et al. (2024), GARAG Cho et al. (2024), and PRCAP Zhong et al. (2023). The dataset is split into 70% for training, 20% for testing, and 10% as a support set. Within the training set, samples are randomly grouped into triplets (anchor, positive, negative), where the anchor and positive belong to the same class, and the negative belongs to a different class.
Table 9: Statistical data and format of the responses.
Dataset Average Word Count of Response An Example of Response
NQ Poisoned Response: 7 Correct Response: 12 Question: where is the food stored in a yam plant? Poisoned Response: In the leaves. Correct Response: In the tuber.
HotpotQA Poisoned Response: 8 Correct Response: 11 Question: Which actor starred in Assignment to Kill and passed away in 2000? Poisoned Response: Patrick O’Neal. Correct Response: John Gielgud.
MS-MARCO Poisoned Response: 16 Correct Response: 24 Question: what is hardie plank? Poisoned Response: Hardie plank is a wood flooring option that is used for a variety of home styles. Correct Response: Hardie Plank is a brand of fiber cement siding.
A.2 Poisoned Texts Generation
To ensure that the retrieved poisoned texts successfully achieve the poisoning effect, we employ three existing methods PoisonedRAG Zou et al. (2024), GARAG Cho et al. (2024), and PRCAP Zhong et al. (2023) to generate the poisoned texts. In the PoisonedRAG Zou et al. (2024) method, the attacker first selects a target question along with its corresponding incorrect answer. The attacker then optimizes the design of the poisoned text to ensure that it meets two key criteria: (1) retrievability by the retriever and (2) effectiveness in misleading the language model to generate the incorrect answer. GARAG Cho et al. (2024) is a novel adversarial attack algorithm that generates adversarial documents by subtly perturbing clean ones while preserving answer tokens. Through iterative crossover, mutation, and selection, it optimizes the documents to maximize adversarial effectiveness within the defined search space. PRCAP Zhong et al. (2023) is a gradient-based method, which starts from a natural-language passage and iteratively perturbs it in the discrete token space to maximize its similarity to a set of training queries.
It’s worth noting that the generation method for poisoned texts can be any approach that successfully achieves the poisoning effect. Once the activations of both correct and poisoned responses are obtained, we preprocess and use them for training and testing the RevPRAG model.
A.3 Prompt
The following is the system prompt for RAG, instructing an LLM to produce a response based on the provided context:
A.4 Environment
We conduct experiments on a server with 64 AMD EPYC 9654 CPUs (64 logical cores enabled) at 2.40–3.70 GHz, 512 GB of DDR5 RAM (assumed based on high-core-count server standards), and four NVIDIA RTX A6000 GPUs, each with 48 GB GDDR6 memory.